Scampini Technologies is expected to generate $50 million in free cash flow next year, and FCF is expected to grow at a constant rate of 5% per year indefinitely. Scampini has no debt or preferred stock, and its WACC is 13%. If Scampini has 65 million shares of stock outstanding, what is the stock's value per share? Round your answer to two decimal places.
Answer #1

Current total value=FCF for next year/(WACC-Growth rate)

=50/(0.13-0.05)=$625million
Hence stock's value per share=(Current total value/Shares of stock oustanding)

=(625/65)

which is equal to

=$9.62(Approx).